A ferry boat has internal volume 1 m^3 and weight 50 kg

If a leak develops in the bottom and water starts coming in, what fraction of the boat's volume will be filled with water starts coming in from the sides? 

Grade:12

1 Answers

India VK
10 years ago

The entire boat is submerged when the water starts coming in from the sides. So volume of water displaced = 1 m^3

Upthrust = wight of water displaced = 1000 * g 

Also suppose v is the vol. of water is inside the boat at this instant.

weight of boat plus weigt of water inside the boat = (50 + v*1000) * g

This should be equal to upthrust = 1000g

so v*1000 = 950 or v = 0.95